# Anomaly Flow: Shi-Type Estimates and Long-time Existence

### v1

@@ 1 Introduction
Suppose that there exists a positive constant $C_{0}$ such that
@@ 3 Integral Shi-Type Estimates
Our objective in this section is to obtain uniform $L^{\infty}$ estimates on covariant derivatives of the curvature and torsion along the anomaly flow under the assumption that the lower derivatives are also uniformly bounded.
@@ 3.1 Starting Assumptions
We note that the first assumption (3.1) follows from the bounds on curvature and torsion from (3.3). In essence, these bounds and (2.2) imply that $|\partial_{t}g|$ is bounded and so the evolving metric $g$ is uniformly bounded above and below by the initial metric. We write (3.1) as a separate assumption since we will derive results based on the constant $B$.
By the same token, (3.3) also implies that the volume $\int_{X}1$ is uniformly bounded along the flow. $\blacktriangle$
@@
As described previously, we do this by first obtaining $L^{2p}$ bounds on $|D^{k}\text{{Rm}}|$ and $|D^{k+1}T|$ under appropriate conditions on $\alpha^{\prime}$. This will be done in §3.2. We then take it one step further in §3.3 and obtain $L^{2p}$ bounds on $|D^{k+1}\text{{Rm}}|$ and $|D^{k+2}T|$. Finally, we apply the Sobolev embedding theorem and an induction argument to get bounds on the higher order derivatives.
@@ 3.2.2 $L^{2p}$ Estimates
where the generic constant $C$ MAY now also depend on $p$. The second term in the first inequality comes from the evolving volume form and that $|\partial_{t}g|$ is assumed to be bounded along the flow.
@@
Using uniform boundedness of the volume, we can retrieve the $L^{2p}$ bounds for $31\leq p<$ 3 via Hölder’s inequality.
@@ 3.4 Sobolev Embedding and Induction
Recall also that $a_{0}$ is a predetermined constant from our expressions (2.5) and (2.8) for $\partial_{t}\text{{Rm}}$ and $\partial_{t}T$ inherent to the anomaly flow.
@@
Suppose this holds for some $p$ with $62p>n=$ 6. The Sobolev embedding theorem on $D^{k}\text{{Rm}}$ and $D^{k+1}T$ then provides $L^{\infty}$ bounds on $D^{k}\text{{Rm}}$ and $D^{k+1}T$ uniform in $t$ along the flow. (See the paragraphs following Lemma 14.3 and also Lemma 14.4 of [Ham82] for more details). That is, there exists some $C_{k}$ such that
@@
Importantly, the condition (3.66) does not depend on $k$ and so we can induct on $k$ to obtain $L^{\infty}$ bounds for $D^{q}\text{{Rm}}$ and $D^{q+1}T$ for each $q\geq 3$. We have thus shown
@@
$\alpha^{\prime}<\frac{1}{16a_{0}B^{2}C_{0}},$ (3.69)
@@ 5 Long-time Existence
Now that we have established the relevant $L^{\infty}$ bounds for the covariant derivatives and torsion, we can appeal to the argument in [PPZ18b] to obtain long-time existence of the anomaly flow. For completeness, we outline the argument below.
@@ 5.1 $C^{0}$- and $C^{1}$-Bounds
Suppose that there exists a positive constant $C_{0}$ such that
@@
along the anomaly flow on $[0,\tau)$. Recall that by Remark 3.1 the dilaton bound
$B^{-1}\leq\Big{(}\frac{1}{2\|\Omega\|_{\omega}}\Big{)}\leq B$ (5.2)
is implied by the existence of this $C_{0}$. The metrics $g$ are uniformly equivalent close to time $\tau$, and so if $\widehat{g}$ is a fixed reference metric, the endomorphism
@@
Since Rm is bounded, we have a $C^{1}$ bound on the endomorphism $h$.

### v2

@@ 1 Introduction
Suppose that there exist positive constants $B$, $C_{0}$ such that
$B^{-1}\leq\Big{(}\frac{1}{2\|\Omega\|_{\omega}}\Big{)}\leq B,$ (1.7)
@@ 3 Integral Shi-Type Estimates
Our objective in this section is to obtain uniform $L^{\infty}$-estimates on covariant derivatives of the curvature and torsion along the anomaly flow under the assumption that the lower derivatives are also uniformly bounded.
@@ 3.1 Starting Assumptions
We note that the first assumption (3.1) and the third assumption (3.3) in conjunction with (2.2) imply that $|\partial_{t}g|$ is uniformly bounded and so the evolving metric $g$ is uniformly bounded above and below by the initial metric. As such, the volume $\int_{X}1$ is uniformly bounded along the flow. $\blacktriangle$
@@
As described previously, we do this by first obtaining $L^{2p}$-bounds on $|D^{k}\text{{Rm}}|$ and $|D^{k+1}T|$ under appropriate conditions on $\alpha^{\prime}$. This will be done in §3.2. We then take it one step further in §3.3 and obtain $L^{2p}$-bounds on $|D^{k+1}\text{{Rm}}|$ and $|D^{k+2}T|$. Finally, we apply the Sobolev embedding theorem and an induction argument to get bounds on the higher order derivatives.
@@ 3.2.2 $L^{2p}$-Estimates
where the generic constant $C$ MAY now also depend on $p$. The second term in the first inequality comes from the evolving volume form and that $|\partial_{t}g|$ is bounded along the flow.
@@
Using uniform boundedness of the volume, we can retrieve the $L^{2p}$-bounds for $31\leq p<$ 3 via Hölder’s inequality.
@@ 3.4 Sobolev Embedding and Induction
Recall also that $a_{0}$ is a predetermined constant from our expressions (2.5) and (2.8) for $\partial_{t}\text{{Rm}}$ and $\partial_{t}T$ that is inherent to the anomaly flow.
@@
Suppose this holds for some $p$ with $62p>n=$ 6. The Sobolev embedding theorem on $D^{k}\text{{Rm}}$ and $D^{k+1}T$ then provides $L^{\infty}$-bounds on $D^{k}\text{{Rm}}$ and $D^{k+1}T$ uniform in $t$ along the flow. (See the paragraphs following Lemma 14.3 and also Lemma 14.4 of [Ham82] for more details). That is, there exists some $C_{k}$ such that
@@
Importantly, the condition (3.66) does not depend on $k$ and so we can induct on $k$ to obtain $L^{\infty}$-bounds for $D^{q}\text{{Rm}}$ and $D^{q+1}T$ for each $q\geq 3$. We have thus shown
@@
$\alpha^{\prime}<\frac{1}{14a_{0}B^{2}C_{0}},$ (3.69)
@@ 5 Long-time Existence
Now that we have established the relevant $L^{\infty}$-bounds for the covariant derivatives and torsion, we can appeal to the argument in [PPZ18b] to obtain long-time existence of the anomaly flow. For completeness, we outline the argument below.
@@ 5.1 $C^{0}$- and $C^{1}$-Bounds
Suppose that there exist positive constants $B$, $C_{0}$ such that
$B^{-1}\leq\Big{(}\frac{1}{2\|\Omega\|_{\omega}}\Big{)}\leq B,$ (5.1)
@@
along the anomaly flow on $[0,\tau)$. The metrics $g$ are uniformly equivalent close to time $\tau$, and so if $\widehat{g}$ is a fixed reference metric, the endomorphism
@@
Since Rm is bounded, we have a $C^{1}$-bound on the endomorphism $h$.